# Gradients: When Markets Meet Fine-tuning - A Distributed Approach to Model Optimisation

### v1

@@
Abstract
Foundation model fine-tuning faces a fundamental challenge: existing AutoML platforms rely on single optimisation strategies that explore only a fraction of viable hyperparameter configurations. In this white paper, We introduce Gradients, a decentralised AutoML platform that transforms hyperparameter optimisation into a competitive marketplace where independent miners compete to discover optimal configurations. Economic incentives align individual exploration with collective optimisation goals, driving systematic investigation of hyperparameter regions that centralised methods miss. We evaluate our approach across 180 controlled experiments spanning diverse model architectures (70M to 70B parameters) and task types. Gradients achieves an 82.8% win rate against HuggingFace AutoTrain and 100% against TogetherAI, Databricks, and Google Cloud, with mean improvements of 11.8% and 42.1% respectively. Complex reasoning and retrieval tasks show particularly strong gains of 30-40%, whilst diffusion models achieve 23.4% improvements for person-specific generation. These results demonstrate that competitive, economically-driven approaches can systematically discover superior configurations that centralised AutoML consistently miss.
@@ Abstract
Figure 1: Statistical analysis by task type. Four-panel analysis: box plots show improvement distributions, bar charts display win rates and mean improvements with standard errors, summary table provides exact values.
1 The open challenges in model fine-tuning.
Foundation model fine-tuning involves optimisation across multiple interdependent dimensions: task type, model architecture, dataset size, and time constraints [41]. A 70M parameter model fine-tuned for mathematical reasoning with limited data requires different strategies than a 70B model trained for translation tasks with extensive datasets. Time constraints further complicate the picture—optimal configurations for 2-hour training windows differ from those for 8-hour allocations. Current AutoML platforms [37, 13] produce results across these diverse scenarios, yet our empirical evaluation across 180 experiments reveals systematic performance gaps, suggesting that task-specific optimisation opportunities remain undiscovered.
The challenge lies in creating systems that can adapt optimisation strategies to these multifaceted problem characteristics. Economic incentives offer a solution [11, 8]: when optimisation success directly determines rewards, agents develop approaches tailored to the specific combination of task type, model scale, dataset properties, and resource constraints.
Such incentive alignment naturally leads to decentralised approaches [24, 17]. By distributing the exploration task across multiple independent agents, each with unique strategies and computational resources, decentralised systems can simultaneously investigate diverse approaches that would otherwise remain unexplored. This parallel search capability scales with the number of participants, creating an explorative capacity that exceeds what centralised systems can achieve with equivalent resources. However, mere distribution of computation is insufficient; the key innovation lies in establishing economic incentives that align individual agent behaviour with collective optimisation goals. When agent rewards are directly tied to discovering superior configurations, the system naturally evolves toward increasingly sophisticated optimisation strategies [20]. This evolutionary pressure creates a form of collective intelligence that continuously adapts to the specific characteristics of each fine-tuning task, rather than applying standardised approaches. The theoretical foundation of this approach combines distributed optimisation [23], empirical risk minimisation, and tournament theory [26], establishing a framework where properly structured competition can efficiently navigate complex optimisation landscapes that resist traditional methods.
In this paper, we present Gradients, a decentralised AutoML platform built on the Bittensor network [32] that implements these theoretical principles to address the fine-tuning challenges of foundation models. We make several contributions: First, we formulate a mathematical scoring framework that effectively ranks model submissions based on their performance across both test and synthetic evaluation datasets, with appropriate mechanisms to prevent overfitting and ensure fair competition. Second, we develop an economic incentive structure that aligns individual miner rewards with overall optimisation quality, creating a self-reinforcing cycle of innovation in hyperparameter exploration. Third, we present empirical results from extensive comparative testing against leading centralised AutoML platforms, demonstrating consistent and significant performance improvements across diverse model types and application domains. Fourth, we analyse the dynamics of the system, showing how the competitive marketplace accelerates the adoption of novel optimisation techniques and enables more efficient allocation of computational resources compared to conventional approaches. Through these contributions, we demonstrate the practical effectiveness of decentralised, economically-driven approaches to hyperparameter optimisation and provide empirical evidence for their advantages, with particular application to the fine-tuning of Large Language Models and diffusion models.
2 Related Work
@@
Hyperparameter Optimisation and AutoML. The field of automated hyperparameter optimisation has progressed from simple grid and random search methods to sophisticated multi-fidelity approaches. Bayesian optimisation frameworks like Hyperband [23] and BOHB [12] combine principled uncertainty estimation with early stopping mechanisms to efficiently navigate large hyperparameter spaces. These methods achieve computational efficiency through bandit-based resource allocation, but remain fundamentally centralised in their coordination mechanisms. Recent AutoML platforms including AutoKeras [21] and Auto-Sklearn [13] have democratised access to these techniques, yet they still rely on single optimisation strategies that may miss task-specific optima. HuggingFace AutoTrain [37] represents the current state-of-the-art in no-code model fine-tuning, providing automated hyperparameter optimisation across multiple modalities but using centralised search strategies.
@@
Distributed and Competitive Optimisation. Population-based training (PBT) [20] introduced competitive dynamics to neural network optimisation, where populations of models compete and evolve hyperparameters during training. This approach demonstrates how competition discovers superior configurations through exploitation and exploration mechanisms reminiscent of evolutionary algorithms [10]. Parallel to this, federated learning [24] established principles for coordinating distributed model training without centralising data, achieving communication efficiency through local computation and periodic aggregation. However, these approaches lack economic incentive mechanisms to align individual participant behaviour with collective optimisation goals.
@@
Parameter-Efficient Fine-tuning. The emergence of foundation models has necessitated new approaches to efficient adaptation. LoRA [19] and its variants inject trainable low-rank matrices into pre-trained transformers, reducing trainable parameters by orders of magnitude whilst preserving performance. AdaLoRA [42] extends this approach with adaptive rank allocation, whilst QLoRA [7] combines quantisation with low-rank adaptation for memory efficiency. These parameter-efficient methods are widely used by AutoML LLM and diffusion training platforms [9].
@@
Economic Mechanisms in Machine Learning. Recent work explores how economic incentives coordinate machine learning systems. Differentiable economics [11] employs neural networks to automatically design optimal auction mechanisms, demonstrating how learning systems embed economic principles. Blockchain-based coordination mechanisms [17, 34] show promise for incentivising distributed computation, whilst tournament theory provides mathematical foundations for competition-based optimisation [26]. Incentive mechanism design for distributed coded machine learning [8] establishes optimal reward structures for heterogeneous computational contributors. However, researchers have largely left unexplored the application of these economic principles to automated hyperparameter optimisation.
@@
Decentralised Machine Learning Networks. Emerging platforms like Bittensor [32] demonstrate how cryptoeconomic incentives coordinate distributed machine learning tasks. These networks employ token-based rewards to align individual computational contributions with network-wide objectives, creating sustainable ecosystems for collaborative AI development. The Yuma Consensus mechanism [27] provides robust peer-ranking systems resistant to collusion, whilst maintaining decentralised coordination. Our work builds on these foundations by specifically applying economic coordination mechanisms to the hyperparameter optimisation problem, demonstrating how competitive markets discover superior model configurations.
@@
Foundation Model Fine-tuning and Evaluation. Recent advances in fine-tuning large language models focus on instruction following [29], preference optimisation [31], and synthetic data generation [2]. For diffusion models, techniques like DreamBooth [33] enable subject-driven generation through personalised fine-tuning. Community-driven platforms like CivitAI [38] have democratised access to fine-tuned diffusion models, though require an understanding of the diffusion training process or the use of default parameters in order to use. Our work addresses this gap by providing economic incentives for discovering optimal fine-tuning configurations across diverse model types and tasks without the need for users of the platform to have a deep understanding of the fine-tuning process.
3 Gradients
@@ 3.1 System Architecture
3.2 Decentralised Task Distribution
@@
$P(\text{TaskType}=t)=\begin{cases}\rho_{\text{instruct}}&\text{if }t=\text{InstructTextTask}\\ \rho_{\text{dpo}}&\text{if }t=\text{DPOTask}\\ \rho_{\text{grpo}}&\text{if }t=\text{GRPOTask}\\ \rho_{\text{image}}&\text{if }t=\text{ImageTask}\end{cases}$ (1)
@@
For each task, the system selects a pool of miners through a weighted probability mechanism designed to balance exploration and exploitation. Given a set of available miners =$\mathcal{M}=\{m_{1},m_{2},...,m_{n}\}$, selection weights are assigned as $w_{i}=\alpha$ if miner $m_{i}$ has not participated today, or $w_{i}=\max(s_{i},\gamma)$ otherwise, where $\alpha$ is the default score for first daily participation (set to 2.0), $s_{i}$ is miner $m_{i}$’s average quality score from previous tasks, and $\gamma$ is the minimum score threshold (0.01). Miners are then sorted by weight and assigned position-based selection probabilities:
$p_{i}=\lambda-(i-1)\cdot\frac{\lambda-1}{|\mathcal{M}|}$ (2)
where $p_{i}$ is the relative probability of selecting the $i$-th ranked miner, $\lambda$ is the top miner chance multiplier (3.0), and $|\mathcal{M}|$ is the total number of available miners. This approach preferentially selects high-performing miners while ensuring all miners maintain participation opportunities, preventing monopolisation by established participants.
3.3 Dataset Preparation Pipeline
Effective dataset preparation ensures fair evaluation and meaningful comparison of fine-tuning approaches. For each task, the system partitions the available data into three distinct sets: $D=D_{\text{train}}\cup D_{\text{test}}\cup D_{\text{synth}}$ where $D_{\text{train}}$ is provided to miners for model fine-tuning, $D_{\text{test}}$ is used for primary performance evaluation with size $|D_{\text{test}}|=\min(|D|\cdot\rho_{\text{test}},\kappa_{\text{test}})$, and $D_{\text{synth}}$ is algorithmically generated to test generalization with size $|D_{\text{synth}}|=\min(|D|\cdot\rho_{\text{synth}},\kappa_{\text{synth}})$, using parameters $\rho_{\text{test}}=0.1$, $\rho_{\text{synth}}=1.0$, $\kappa_{\text{test}}=1,000$, and $\kappa_{\text{synth}}=300$. Synthetic data generation serves multiple purposes: it provides an additional evaluation benchmark, tests for generalisation capabilities, and mitigates overfitting to the test set. For language models, using a controlled process with temperature $\tau=0.6$, we generate synthetic data that preserves schema requirements while introducing semantic diversity through the process $x^{\prime}_{i}=G(x_{i},\tau)$ for all $x_{i}\in D_{\text{sample}}$, where $G$ is a high-quality foundation model generating function. For DPO tasks, this includes generating both chosen and rejected responses using models of different capabilities. For diffusion models, using temperature $\tau=0.4$, the system generates new image-text pairs $(I_{i},P_{i})=G_{\text{image}}(S_{i},R_{i})$ with controlled style variations and prompt diversity, where $S_{i}$ is a style specification and $R_{i}$ is a target resolution. Image resolutions are standardised to ensure fair comparison, with dimensions constrained to be divisible by 64.
3.3.1 Task Time Allocation
@@
$\Delta t\in\begin{cases}[3,6]\text{ hours}&\text{if }|D|\in[10000,25000]\\ [4,8]\text{ hours}&\text{if }|D|\in[25000,50000]\\ [5,9]\text{ hours}&\text{if }|D|\in[50000,100000]\\ [7,10]\text{ hours}&\text{if }|D|\in[100000,500000]\end{cases}$ (4)
@@
3.4 Multi-Level Evaluation Framework
To ensure fair comparison and prevent gaming, all models are evaluated in isolated Docker environments with standardised GPU configurations. This creates a controlled evaluation sandbox where performance differences reflect genuine optimisation quality rather than environmental variations. For language models, the evaluation process computes both test and synthetic losses:
$L_{\text{test}}(m_{i})$ $=\frac{1}{|D_{\text{test}}|}\sum_{x\in D_{\text{test}}}\ell(M_{m_{i}},x)$ (5)
$L_{\text{synth}}(m_{i})$ $=\frac{1}{|D_{\text{synth}}|}\sum_{x\in D_{\text{synth}}}\ell(M_{m_{i}},x)$ (6)
where $M_{m_{i}}$ is the model submitted by miner $m_{i}$ and $\ell$ is an appropriate loss function. For diffusion models, evaluation encompasses text-guided and non-text-guided image reconstruction quality:
$L_{\text{text}}(m_{i})$ $=\frac{1}{|D_{\text{test}}|}\sum_{(I,p)\in D_{\text{test}}}\|I-M_{m_{i}}(I_{\text{noisy}},p)\|_{2}^{2}$ (7)
$L_{\text{no-text}}(m_{i})$ $=\frac{1}{|D_{\text{test}}|}\sum_{I\in D_{\text{test}}}\|I-M_{m_{i}}(I_{\text{noisy}},\emptyset)\|_{2}^{2}$ (8)
where $I$ is the original target image, $I_{\text{noisy}}$ is the noise-corrupted input, and $p$ is the text prompt. To prevent overfitting to either evaluation set, performance is measured as a weighted combination $L_{\text{weighted}}(m_{i})=\omega\cdot L_{\text{test}}(m_{i})+(1-\omega)\cdot L_{\text{synth}}(m_{i})$ where $\omega$ is the test score weighting parameter (0.7 for text tasks). For image models, the weighting combines text-guided and non-text-guided metrics: $L_{\text{weighted}}(m_{i})=\delta\cdot L_{\text{text}}(m_{i})+(1-\delta)\cdot L_{\text{no-text}}(m_{i})$ with $\delta=0.25$ emphasising coherence over prompt adherence.
3.4.1 Anti-Gaming Mechanisms
To maintain system integrity, several mechanisms detect and penalise attempts to game the evaluation:
Duplicate Detection:
Submissions with identical loss patterns are identified, and only the earliest submission is credited:
$\text{isDuplicate}(m_{i},m_{j})=\begin{cases}\text{true}&\text{if }|L_{\text{test}}(m_{i})-L_{\text{test}}(m_{j})|<\epsilon\\ &\text{and }|L_{\text{synth}}(m_{i})-L_{\text{synth}}(m_{j})|<\epsilon\\ \text{false}&\text{otherwise}\end{cases}$ (9)
Suspicious Performance Detection:
Models showing anomalous performance disparities between test and synthetic datasets are flagged:
$\text{isSuspicious}(m_{i})=\begin{cases}\text{true}&\text{if }L_{\text{synth}}(m_{i})>L_{\text{test}}(m_{i})+\alpha\cdot\sigma(L_{\text{test}})\\ \text{false}&\text{otherwise}\end{cases}$ (10)
where $\alpha$ is a sensitivity parameter (0.5) and $\sigma(L_{\text{test}})$ is the standard deviation of test losses.
3.5 Miner Scoring

### v2

@@
Abstract.
Current AutoML platforms leave substantial performance untapped. Testing 180 fine-tuning tasks across models from 70M to 70B parameters, we found that HuggingFace AutoTrain, TogetherAI, Databricks, and Google Cloud consistently produce suboptimal configurations. Gradients, built on the Bittensor network, attacks this problem through competition. Independent miners race to find optimal hyperparameters, earning rewards proportional to their models’ performance. This tournament drives exploration of configuration spaces that single-strategy methods never examine. In our experiments, Gradients achieved a 100% win rate against TogetherAI, Databricks, and Google Cloud, and beat HuggingFace AutoTrain in 82.8% of experiments. Mean improvements reached 42.1% against commercial platforms. Retrieval-augmented generation tasks saw 30-40% gains; diffusion models improved 23.4% on person-specific generation. When miners compete for rewards, they develop optimization strategies that centralized approaches overlook. These findings demonstrate that decentralized systems with economic incentives can systematically outperform traditional AutoML, suggesting market dynamics may be key to achieving superior fine-tuning results. Code is available at https://github.com/rayonlabs/G.O.D.
copyright: nonejournalyear: 2025
@@ Abstract.
Figure 1. Performance improvement by task type. Four-panel analysis: box plots show improvement distributions, bar charts display win rates and mean improvements with standard errors, summary table provides exact values.
1. The open challenges in model fine-tuning
Fine-tuning foundation models has become a critical component of modern AI deployment, yet current AutoML platforms consistently produce suboptimal configurations. Our evaluation across 180 fine-tuning tasks reveals that HuggingFace AutoTrain, TogetherAI, Databricks, and Google Cloud leave performance gaps of 20-40% in retrieval and reasoning tasks, suggesting that current approaches leave substantial performance unrealized.
The root cause lies in the nature of the fine-tuning problem itself. Fine-tuning is not a single optimization challenge but thousands of distinct problems (Zhang et al., 2024). A 7B parameter model learning to generate code requires fundamentally different hyperparameters than the same model learning Japanese translation. A 70M model trained on 10,000 examples for 2 hours needs different settings than a 70B model trained on 500,000 examples for 10 hours. Current AutoML platforms (Thakur, 2024; Feurer et al., 2015) address this diversity with single optimization strategies—whether Bayesian optimization, grid search, or evolutionary algorithms. These methods explore only a small fraction of the viable configuration space. They cannot adapt to the specific characteristics of each task, model, and dataset combination. The result is systematic underperformance: acceptable results that miss superior configurations.
Mechanism design and distributed optimization research offer an alternative approach (Dütting et al., 2019; Ding et al., 2020). When solution spaces are vast and optimal strategies unknown, competitive mechanisms often outperform centralized planning (McMahan et al., 2017; Harris and Waggoner, 2019). Consider how Linux development leverages thousands of independent contributors, each pursuing different improvements. Or how Bitcoin mining incentivizes global competition to secure the network. In machine learning specifically, population-based training (Jaderberg et al., 2017) suggests that competition between agents often lead to better solutions than single-agent optimization. These examples share a common principle: properly aligned incentives drive exploration of solution spaces that centralized approaches cannot effectively navigate.
Gradients implements this principle for foundation model fine-tuning. Built on the Bittensor network (Rao et al., 2021), it creates a marketplace where independent miners compete to find optimal configurations. When a fine-tuning task arrives, the system distributes it to multiple miners—typically 10-20 for language models, 15-25 for diffusion models. Each miner independently determines their approach: learning rate schedules, batch sizes, optimization algorithms, LoRA configurations, data augmentation strategies. They train models in parallel, each exploring different regions of the hyperparameter space. The system evaluates all submissions on held-out test data and synthetic benchmarks, rewarding miners proportionally to their models’ performance. This creates direct economic incentives for discovering superior configurations.
The empirical results demonstrate the effectiveness of this approach. Across 180 controlled experiments, Gradients achieved a 100% win rate against TogetherAI, Databricks, and Google Cloud. It outperformed HuggingFace AutoTrain—widely considered the state-of-the-art in automated fine-tuning—in 82.8% of direct comparisons. Mean improvements reached 11.8% against HuggingFace and 42.1% against commercial cloud platforms. Retrieval-augmented generation tasks showed gains of 30-40%, while diffusion models improved 23.4% on person-specific generation tasks. These improvements span models from 70M to 70B parameters and diverse task types including translation, code generation, mathematical reasoning, and image generation.
These findings have implications beyond immediate performance gains. As foundation models grow larger and tasks become more specialized, the configuration space expands exponentially. No single algorithm, regardless of sophistication, can adequately explore these vast spaces. Competitive markets, however, naturally parallelize exploration across diverse strategies. When miners profit from finding better configurations, they develop novel optimization techniques, share successful strategies within their organizations, and continuously refine their approaches. The system’s knowledge grows through evolutionary pressure rather than algorithmic updates. This suggests a fundamental shift in how we approach AutoML: from better algorithms to better incentive structures.
This paper presents the design, implementation, and evaluation of the Gradients system. Section 2 reviews related work in hyperparameter optimization, distributed machine learning, and economic mechanism design. Section 3 describes the Gradients architecture, including task distribution, miner selection, evaluation protocols, and reward mechanisms. Section 4 details our experimental methodology for comparing against existing AutoML platforms. Section 5 presents comprehensive results across model scales and task types. Section 6 discusses why competitive approaches succeed and their implications for future AutoML systems. We conclude by considering the broader potential of economically-driven optimization in machine learning. The theoretical foundation of this approach combines distributed optimization (Li et al., 2017), empirical risk minimization, and tournament theory (Moldovanu and Sela, 2001), establishing a framework where properly structured competition can efficiently navigate complex optimization landscapes that resist traditional methods.
2. Related Work
2.1. Hyperparameter Optimisation and AutoML
@@
The field of automated hyperparameter optimisation has progressed from simple grid and random search methods to sophisticated multi-fidelity approaches. Bayesian optimisation frameworks like Hyperband (Li et al., 2017) and BOHB (Falkner et al., 2018) combine principled uncertainty estimation with early stopping mechanisms to efficiently navigate large hyperparameter spaces. These methods achieve computational efficiency through bandit-based resource allocation, but remain fundamentally centralised in their coordination mechanisms. Recent AutoML platforms including AutoKeras (Jin et al., 2019) and Auto-Sklearn (Feurer et al., 2015) have democratised access to these techniques, yet they still rely on single optimisation strategies that may miss task-specific optima. HuggingFace AutoTrain (Thakur, 2024) represents the current state-of-the-art in no-code model fine-tuning, providing automated hyperparameter optimisation across multiple modalities but using centralised search strategies.
2.2. Distributed and Competitive Optimisation
@@
Population-based training (PBT) (Jaderberg et al., 2017) introduced competitive dynamics to neural network optimisation, where populations of models compete and evolve hyperparameters during training. This approach demonstrates how competition discovers superior configurations through exploitation and exploration mechanisms reminiscent of evolutionary algorithms (Dushatskiy et al., 2023). Parallel to this, federated learning (McMahan et al., 2017) established principles for coordinating distributed model training without centralising data, achieving communication efficiency through local computation and periodic aggregation. However, these approaches lack economic incentive mechanisms to align individual participant behaviour with collective optimisation goals.
2.3. Parameter-Efficient Fine-tuning
@@
The emergence of foundation models has necessitated new approaches to efficient adaptation. LoRA (Hu et al., 2022) and its variants inject trainable low-rank matrices into pre-trained transformers, reducing trainable parameters by orders of magnitude whilst preserving performance. AdaLoRA (Zhang et al., 2023) extends this approach with adaptive rank allocation, whilst QLoRA (Dettmers et al., 2023) combines quantisation with low-rank adaptation for memory efficiency. These parameter-efficient methods are widely used by AutoML LLM and diffusion training platforms (Ding et al., 2024).
2.4. Economic Mechanisms in Machine Learning
@@
Recent work explores how economic incentives coordinate machine learning systems. Differentiable economics (Dütting et al., 2019) employs neural networks to automatically design optimal auction mechanisms, demonstrating how learning systems embed economic principles. Blockchain-based coordination mechanisms (Harris and Waggoner, 2019; Shayan et al., 2021) show promise for incentivising distributed computation, whilst tournament theory provides mathematical foundations for competition-based optimisation (Moldovanu and Sela, 2001). Incentive mechanism design for distributed coded machine learning (Ding et al., 2020) establishes optimal reward structures for heterogeneous computational contributors. However, researchers have largely left unexplored the application of these economic principles to automated hyperparameter optimisation.
2.5. Decentralised Machine Learning Networks
@@
Emerging platforms like Bittensor (Rao et al., 2021) demonstrate how cryptoeconomic incentives coordinate distributed machine learning tasks. These networks employ token-based rewards to align individual computational contributions with network-wide objectives, creating sustainable ecosystems for collaborative AI development. The Yuma Consensus mechanism (Opentensor Foundation, 2025) provides robust peer-ranking systems resistant to collusion, whilst maintaining decentralised coordination. Our work builds on these foundations by specifically applying economic coordination mechanisms to the hyperparameter optimisation problem, demonstrating how competitive markets discover superior model configurations.
2.6. Foundation Model Fine-tuning and Evaluation
@@
Recent advances in fine-tuning large language models focus on instruction following (Ouyang et al., 2022), preference optimisation (Rafailov et al., 2024), and synthetic data generation (Chan et al., 2024). For diffusion models, techniques like DreamBooth (Ruiz et al., 2023) enable subject-driven generation through personalised fine-tuning. Community-driven platforms like CivitAI (Wei et al., 2024) have democratised access to fine-tuned diffusion models, though require an understanding of the diffusion training process or the use of default parameters in order to use. Our work addresses this gap by providing economic incentives for discovering optimal fine-tuning configurations across diverse model types and tasks without the need for users of the platform to have a deep understanding of the fine-tuning process.
3. Gradients
@@ 3.1. System Architecture
3.2. Decentralised Task Distribution
@@
[…]
@@
For each task, the system selects a pool of miners through a weighted probability mechanism designed to balance exploration and exploitation. Given a set of available miners $\mathcal{M}=\{m_{1},m_{2},...,m_{n}\}$, selection weights are assigned as $w_{i}=\alpha$ if miner $m_{i}$ has not participated today, or $w_{i}=\max(s_{i},\gamma)$ otherwise, where $\alpha$ is the default score for first daily participation (set to 2.0), $s_{i}$ is miner $m_{i}$’s average quality score from previous tasks, and $\gamma$ is the minimum score threshold (0.01).
Miners are then sorted by weight and assigned position-based selection probabilities:
(2) $p_{i}=\lambda-(i-1)\cdot\frac{\lambda-1}{|\mathcal{M}|}$
where $p_{i}$ is the relative probability of selecting the $i$-th ranked miner, $\lambda$ is the top miner chance multiplier (3.0), and $|\mathcal{M}|$ is the total number of available miners.
This approach preferentially selects high-performing miners while ensuring all miners maintain participation opportunities, preventing monopolisation by established participants.
3.3. Dataset Preparation Pipeline
Effective dataset preparation ensures fair evaluation and meaningful comparison of fine-tuning approaches. For each task, the system partitions the available data into three distinct sets: $D=D_{\text{train}}\cup D_{\text{test}}\cup D_{\text{synth}}$ where $D_{\text{train}}$ is provided to miners for model fine-tuning, $D_{\text{test}}$ is used for primary performance evaluation with size $|D_{\text{test}}|=\min(|D|\cdot\rho_{\text{test}},\kappa_{\text{test}})$, and $D_{\text{synth}}$ is algorithmically generated to test generalization with size $|D_{\text{synth}}|=\min(|D|\cdot\rho_{\text{synth}},\kappa_{\text{synth}})$, using parameters $\rho_{\text{test}}=0.1$, $\rho_{\text{synth}}=1.0$, $\kappa_{\text{test}}=1,000$, and $\kappa_{\text{synth}}=300$.
Synthetic data generation serves multiple purposes: it provides an additional evaluation benchmark, tests for generalisation capabilities, and mitigates overfitting to the test set. For language models, using a controlled process with temperature $\tau=0.6$, we generate synthetic data that preserves schema requirements while introducing semantic diversity through the process $x^{\prime}_{i}=G(x_{i},\tau)$ for all $x_{i}\in D_{\text{sample}}$, where $G$ is a high-quality foundation model generating function. For DPO tasks, this includes generating both chosen and rejected responses using models of different capabilities. For diffusion models, using temperature $\tau=0.4$, the system generates new image-text pairs $(I_{i},P_{i})=G_{\text{image}}(S_{i},R_{i})$ with controlled style variations and prompt diversity, where $S_{i}$ is a style specification and $R_{i}$ is a target resolution. Image resolutions are standardised to ensure fair comparison, with dimensions constrained to be divisible by 64.
3.3.1. Task Time Allocation
@@
[…]
@@
3.4. Multi-Level Evaluation Framework
To ensure fair comparison and prevent gaming, all models are evaluated in isolated Docker environments with standardised GPU configurations. This creates a controlled evaluation sandbox where performance differences reflect genuine optimisation quality rather than environmental variations.
For language models, the evaluation process computes both test and synthetic losses:
(5) $L_{\text{test}}(m_{i})$ $=\frac{1}{|D_{\text{test}}|}\sum_{x\in D_{\text{test}}}\ell(M_{m_{i}},x)$
(6) $L_{\text{synth}}(m_{i})$ $=\frac{1}{|D_{\text{synth}}|}\sum_{x\in D_{\text{synth}}}\ell(M_{m_{i}},x)$
where $M_{m_{i}}$ is the model submitted by miner $m_{i}$ and $\ell$ is an appropriate loss function.
For diffusion models, evaluation encompasses text-guided and non-text-guided image reconstruction quality:
(7) $L_{\text{text}}(m_{i})$ $=\frac{1}{|D_{\text{test}}|}\sum_{(I,p)\in D_{\text{test}}}\|I-M_{m_{i}}(I_{\text{noisy}},p)\|_{2}^{2}$
(8) $L_{\text{no-text}}(m_{i})$ $=\frac{1}{|D_{\text{test}}|}\sum_{I\in D_{\text{test}}}\|I-M_{m_{i}}(I_{\text{noisy}},\emptyset)\|_{2}^{2}$
where $I$ is the original target image, $I_{\text{noisy}}$ is the noise-corrupted input, and $p$ is the text prompt.
To prevent overfitting to either evaluation set, performance is measured as a weighted combination $L_{\text{weighted}}(m_{i})=\omega\cdot L_{\text{test}}(m_{i})+(1-\omega)\cdot L_{\text{synth}}(m_{i})$ where $\omega$ is the test score weighting parameter (0.7 for text tasks). For image models, the weighting combines text-guided and non-text-guided metrics: $L_{\text{weighted}}(m_{i})=\delta\cdot L_{\text{text}}(m_{i})+(1-\delta)\cdot L_{\text{no-text}}(m_{i})$ with $\delta=0.25$ emphasising coherence over prompt adherence.
3.5. Miner Scoring